OpenStudy (kendricklamar2014):

Determine if the number 5.1 * 10^-8 is written in scientific notation. If not, explain. A. Yes; the number is written in scientific notation. B. No; the first factor is not a number between 1 and 10. C. No; it is not written as a number times a power of 10.

OpenStudy (solomonzelman):

what do you think (based on what I said in the prev. question) ?

OpenStudy (kendricklamar2014):

Would the answer be A @SolomonZelman

OpenStudy (solomonzelman):

yes

OpenStudy (anonymous):

should be

OpenStudy (solomonzelman):

remember, it is: \(\large\color{black}{ \displaystyle {\rm C}\times 10^{{\rm N}} }\) where: \(\large\color{black}{ \displaystyle 0>{\rm C}>10 }\) good \(\large\color{black}{ \displaystyle {\rm N}~~~is~~an~~integer! }\) good

OpenStudy (kendricklamar2014):

Yea, thanks you @SolomonZelman and @autumn421-eva

OpenStudy (solomonzelman):

ALL CONDITIONS for it to be a number written in a scientific notation ARE SATISFIED.

OpenStudy (solomonzelman):

yw